# Moyun: A Diffusion-Based Model for Style-Specific Chinese Calligraphy Generation

### v1

@@ IV-B Experiment Setup
For the choice of VAE[16], we used the same pre-trained VAE as in LDM[17]. Specifically, the original image of $256\times 256\times 3$ size is mapped to a latent space of $32\times 32\times 4$ size. We set the number of iterations $N$ for our model to 4, with a hidden layer depth of 512, and the image segmentation patch size to 8. During training, we used a learning rate of $41e-$ 4 and trained on three A100 GPUs with a global batch size of 768. Ultimately, we selected the model at 288,000 steps (19,199 epochs) for subsequent experiment.
IV-C Evaluation Metrics
We evaluated our model from two perspectives: the structure and the style of the generated calligraphy. To assess the basic structural accuracy, we used Tencent’s handwritten OCR service [^2] https://cloud.tencent.com/document/product/866/36212 to recognize the generated calligraphy. Additionally, we used objective parameters such as IOU (Intersection over Union) and PSNR (Peak Signal-to-Noise Ratio) to measure the similarity between the generated calligraphy and the ground truth. A higher similarity indicates a closer match in style and a more reasonable structure. Lastly, we conducted a human evaluation experiment to verify how well our model captures the calligraphy style.
IV-D Experiment Results
Qualitative Evaluation
@@ IV-D Experiment Results
Tencent’s handwritten OCR service was used to recognize the generated calligraphy. The generated calligraphy were provided to the OCR API, which returned the recognized characters. We compared these with the character labels used to guide the generation process. If the recognized character matched the character label, we considered the generated character to have passed the test, indicating good structural integrity.
We randomly selected 60 images from each of the five different fonts for evaluation, considering that the recognition accuracy of handwritten OCR varies across different fonts. The results are shown in Table I. The models and datasets of other works have not been open-sourced, so this experiment was not conducted.
@@
Moyun 0.783 0.583 0.167 0.117 0.03
TABLE I: OCR test result
@@
We generated the same number of images as the entire test set using the prompts from the test set and evaluated the IOU (Intersection over Union). The results showed that our model significantly outperforms other models in terms of IOU. Using the same test set, we also evaluated the PSNR (Peak Signal-to-Noise Ratio) of the model. The results demonstrated that our model has a better performance in PSNR either. These findings are presented in Table II. This indicates that our model is better at fitting the structural integrity of calligraphy characters and more accurately replicating the style of calligraphy.
@@
Zipeng Zhao[13] - 25.3734
Moyun(Ours) 0.810 32.0727
TABLE II: Qualitative measurements
Quantitative Evaluation
We designed a human evaluation experiment by selecting ten generated calligraphy samples and creating ten corresponding questions. Some of these calligraphy was shown in Figure 4. For each question, we provided one generated calligraphy, with four options: one calligraphy from the backbone, and three calligraphy from other calligraphers’ works. Both of them are the same character. The question posed was, “Please select the calligraphy that most closely matches the style of the given one.”
Figure 4: Generated calligraphy. Each column with different labels. The first row shows calligraphy generated by the model which were unseen before. and the second row is ground truth.
The questionnaire was distributed to 9 participants , and the results shows that $53.3$% generated calligraphy was paired with ground truth. This demonstrates that our model is capable of generating calligraphy in the specified style. The models and datasets of other works have not been open-sourced, so they were not included in the questionnaire.
V conclusion

### v2

@@ 4.2. Experiment Setup
In terms of model configuration, we set the number of iterations $N$ for our model to 4, with a hidden layer dimensionality of 512. The image segmentation patch size was set to 8. During training, we used a fixed learning rate of 1e41e-4 and trained the model on three A100 GPUs with a global batch size of 768.
After extensive training, we selected the model checkpoint at 288,000 training steps (approximately 19,199 epochs) for all subsequent experiments.
4.3. Evaluation Metrics
We evaluated the performance of our model from two complementary perspectives: structural fidelity and stylistic authenticity , which are both critical for high-quality calligraphy generation.
To assess the structural accuracy of the generated calligraphy, we employed Tencent’s handwritten Optical Character Recognition (OCR) service [^2] https://cloud.tencent.com/document/product/866/36212 to recognize the content of each generated image. The OCR recognition results were then compared with the ground-truth character labels. A higher recognition accuracy indicates that the generated character retains a structure that is close to the original and remains legible to real-world systems.
In addition to this semantic-level evaluation, we also adopted two widely used objective metrics—Intersection over Union (IoU) and Peak Signal-to-Noise Ratio (PSNR) — to quantitatively measure the pixel-level similarity between the generated images and their corresponding ground truth. These metrics provide an indication of how well the model reconstructs the visual content, reflecting both structural consistency and stylistic resemblance. Higher values on these metrics generally suggest better alignment in terms of layout, stroke formation, and overall appearance.
Beyond automated evaluations, we also conducted out a subjective human assessment to further evaluate the stylistic authenticity and visual quality of the generated calligraphy. Specifically, we invited a group of professional calligraphers and art students to evaluate the outputs in a blind test setting. This qualitative analysis complements the quantitative results by capturing aspects that are difficult to measure with objective metrics, such as artistic expression and resemblance to traditional calligraphic styles.
4.4. Experiment Results
@@
Qualitative Evaluation We employed Tencent’s handwritten OCR service to evaluate the structural integrity of the generated calligraphy. Specifically, each generated image was submitted to the OCR API, which returned the recognized character. This result was then compared with the original character label used during generation. If the two matched, the sample was considered to have passed the structural test, indicating that the generated character retained a legible and recognizable form.
The OCR recognition results are summarized in Table 1, which reports the number of successfully recognized characters across different font styles. For comparison, the table also includes the results obtained on the ground truth (GT) images, as well as the total number of test samples per font category. Note that the models and datasets from previous works have not been made publicly available; therefore, we were unable to conduct a direct comparison with those methods.
@@
Moyun(Ours) 219 177 25 17 1
GT 267 240 51 59 5
Total 351 421 347 109 67
Table 1. OCR test result
@@
We generated the same number of images as the entire test set using the prompts from the test set and evaluated the IOU (Intersection over Union). The results showed that our model significantly outperforms other models in terms of IOU. Using the same test set, we also evaluated the PSNR (Peak Signal-to-Noise Ratio) of the model. The results demonstrated that our model has a better performance in PSNR either. These findings are presented in Table 2. This indicates that our model is better at fitting the structural integrity of calligraphy characters and more accurately replicating the style of calligraphy.
@@
Zipeng Zhao(zhou_end--end_2021) - 25.3734
Moyun(Ours) 0.260 32.0727
Table 2. Qualitative measurements
Quantitative Evaluation To rigorously assess the performance and stylistic fidelity of our generated calligraphy, we designed a human evaluation questionnaire. This questionnaire aimed to gauge how closely our model’s outputs resemble authentic works from specific calligraphers, thereby providing an objective measure of the model’s ability to capture and reproduce nuanced artistic styles.
We selected ten representative calligraphy samples generated by our model, each corresponding to a distinct character. These samples were chosen based on their visual quality and stylistic diversity, ensuring a comprehensive evaluation across various aspects of calligraphic artistry. Some of these calligraphy was shown in Figure 5.
In our questionnaire, each question is based on a calligraphy image generated by our model, which serves as the prompt. Below the prompt, four options are presented, each showing a real-world calligraphy sample of the same character written by a different renowned calligrapher. Among these four options, only one corresponds to the same TripleLabel used to generate the prompt image—this option is referred to as the ”Backbone” option. Participants are asked to select the sample that most closely matches the style of the generated image. This design allows us to evaluate whether the generated results are perceived as stylistically consistent with the intended real calligrapher.
Figure 5. Generated calligraphy. Each column with different labels. The first row shows calligraphy generated by the model which were unseen before. and the second row is ground truth.
The questionnaire was administered to 9 participants, including both calligraphy enthusiasts and individuals with general interest in Chinese art, to evaluate the stylistic fidelity of our generated calligraphy. Each participant was presented with 10 questions, resulting in a total of 90 responses. The results show that 53.353.3% of the generated calligraphy samples were correctly paired with their corresponding ground-truth style—i.e., participants selected the ”Backbone” option (the real calligrapher matching the TripleLabel) as the most stylistically similar sample.
This accuracy rate, significantly above the random guessing baseline of 2525% (chance level for four choices), indicates that our model successfully captures and reproduces key stylistic features of the target calligraphers. It suggests that the generated characters are not only visually plausible but also semantically aligned with the intended artistic style, to the extent that human observers can reliably detect the stylistic correspondence despite the subtle differences inherent in calligraphic expression.
Furthermore, the result supports the effectiveness of our approach in leveraging the TripleLabel to condition the generation process, enabling fine-grained control over stylistic output. While there remains room for improvement—particularly in handling highly nuanced or idiosyncratic styles—this human evaluation provides strong qualitative evidence that our model generates perceptually coherent and stylistically consistent calligraphy, bridging the gap between synthetic generation and authentic artistic expression. The models and datasets of other works have not been open-sourced, so they were not included in the questionnaire.
5. Conclusions
@@
Acknowledgements.
This work was supported by the Shanghai Municipal Science and Technology Commission (24511106802), and the computation is performed in ECNU Multifunctional Platform for Innovation(001).